OpenStudy (anonymous):

while adding the first few continous natural numbers, a candidate missed one of the number and wrote the answer as 177. what was the number missed ?

ganeshie8 (ganeshie8):

find the first triangular number after 177, and subtract 177 from that

OpenStudy (confusionist):

Begin with an equation(let K be the number for now): n(n+1)/2-k=177 13=k I can give you more information if you need it, but that seems like a pain.

ganeshie8 (ganeshie8):

^^ that works too

OpenStudy (anonymous):

okeyplezsend more details @confusionist

OpenStudy (inkyvoyd):

HINT: the sum of the first 19 numbers =19(19+1)/2=190

OpenStudy (inkyvoyd):

well, excluding 0

OpenStudy (confusionist):

I assume you just want to know how to do the equation: n(n+1)/2-k=177 now? Is that what you mean by more information?

OpenStudy (anonymous):

yesss

OpenStudy (confusionist):

@inkyvoyd I'm having trouble doing this without my tablet pen. Mind if you go ahead and solve this? Sorry, I'm glitchy today. x_x

OpenStudy (inkyvoyd):

@suman2335 , as @ganeshie8 said, you want to find the largest smallest value of n that is greater than 177. after trial and error, you will find that for n=19, you get a value of 190. In other words, the sum of 1+2+3+4+...+19=190. Thus, you know that 190-k=177. Solving for k, you get k=13.

OpenStudy (anonymous):

thanks